INAF queue observing, 2016 November 7 - 8

Observer: F. Cusano, A. Rossi
Telescope Operator: D. Gonzalez
Support Astronomer: D. Thompson on call
Instrument: LUCI

Summary

At the sunset the sky is clear
LUCI2 still not available
We use LUCI1.

Night duration (12\xB0 el twilight) 18:30-5:30 ( 01:30-12.30 UT) = 11h 00m

Observing time: 8.3h (76%)

Weather time loss: 2h (18%)

technical time loss 0.66h (6%)

Data Summary

Project_RUN INST Objects Filter ExpTime Notes
Ricci B1 LUCI1 B1 G210 J 1.5 completed B1; RICCI COMPLETED
Antoniucci LUCI1 V2492_Cyg zJ+HK 19m=0.3h missing imaging, guide star is wrong
Garofalo LBC stream Tr-Pisc B,V 25min sequence 6-7-8-9-10 are done
Annibali R3 LBC UGC1281 g,r 25m OB3+calibration OB. R3 completed
Annibali R4 LBC KK16 g,r 30m OB2 only half. OB3 TBD
Calibration LUCI darks   300,200,80 for Iovino & Ricci, Antoniucci

Details

0:40 starting LUCI skyflats in Ks
  • flat not good because the guider was inside the patrol field; persistence we try to flush the detector executing darks

1:32 slewing to telluric Ricci B1, FRicci_1617_20_B1_LUCI1_tell.xml
  • spectra luci1.20161108.0034-0035
  • seeing 1" on guider at 1:39

1:44 slewing to target Ricci B1, FRicci_1617_20_B1_LUCI1.xml
  • spectra luci1.20161108.0040-0057
  • seeing 1.1" at start
  • ~02:15-02:25 seeing ~1.5-2
  • seeing 1.2" at 2:27
  • 02:30 send OB2
  • seeing 1" at 2:33
  • seeing 1.2" at 3:00
  • 03:05 send OB3
  • loss 5 min due to problem with grating G210 central wavelength.
    Super hero Andrea saved the situation moving to mirror and back to G210 1.15 again
  • seeing 1.2" at 3:13
  • seeing 1.2" at 3:20 seeing 1.5"
  • seeing 1.2-2.0" at 3:30
  • Even if seeing >1.2" sometimes, we consider Ricci B1 completed given that in September other exposures where taken. <br\> We can clearly see the emission lines in raw files

03:46 Slew to Antoniucci_6 V2492_Cyg
  • Start telluric
  • luci1.20161108.0061-0064
  • 04:00 start longslit for Antoniucci_6 V2492_Cyg
  • seeing 1.2" at 04:16
  • spectra luci1.20161108.0069-0072 problem with offsets or acquisition, not good!
  • 04:44 Start script again
  • seeing 1.5" at 04:45
  • spectra luci1.20161108.0077-0080 zJ spectra These are good !
  • seeing 1.4" at 04:51
  • spectra luci1.20161108.0081-0084 HK spectra These are good !
  • seeing 1.6" at 05:10
  • 05:16 loss 5min problem with mask. guiding star for imaging is wrong

05:23 Slew to Piconcelli A
  • loss 15min problem with GUI it is not sending the preset to telescope

5:40 we can not solve the problem with LUCI we switch to LBC.
In the mean time Dave is working to solve the problem, in case we still need LUCI

6:03 slewing to copointing of Garofalo, TP_A_point.ob (starting with bad seeing program)
  • dofpia, /X2
  • seeing 1.1" on align images

06:16 science of Garogalo, executing sequence TP_A_6.ob, TP_B_6.ob, TP_C_6.ob, TP_D_6.ob, TP_E_6.ob, FIELD_6.ob
  • seeing 1.5" on images at 6:20
  • seeing 1.2" on images at 6:24
  • seeing 1.1" on images at 6:30
  • 6:37 seeing is stable below 1.2" we move to the higher priority program of Annibali

06:38 slew to Annibali R3
  • dofpia
  • 6:43 executing UGC1281_3.ob
  • seeing 1.1" at 6:50
  • seeing 1" at 6:54
  • added 2 exposures more + calibration OB

07:18 slew to Annibali R4
  • dofpia
  • 07:25 start OB1
  • seeing 1.1" at 07:35
  • seeing 1.1" at 7:51
  • 07:55 start OB2
  • seeing 1.3" at 8:02
  • dofpia
  • seeing 1.6" at 8:10
  • 8:14 OB2 stop. Only half executed, better to repeat

08:16 DIMM seeing 2.2 Back to Garofalo, sequence 7
  • preset problem loss 15min
  • 08:32 dofpia
  • 08:38 start sequence 7: TP_A_7.ob, TP_B_7.ob, TP_C_7.ob, TP_D_7.ob, TP_E_7.ob, FIELD_7.ob
  • seeing 2.5" at 08:40
  • DIMM seeing 3 at 08:50
  • seeing 2" at 08:58 on image

09:02 Garofalo, sequence 8: TP_A_8.ob, TP_B_8.ob, TP_C_8.ob, TP_D_8.ob, TP_E_8.ob, FIELD_8.ob
  • 09:01 seeing 1.4"
  • 09:05 seeing 1.8"
  • 09:15 seeing 2.5"
  • dofpia
  • 09:22 dofpia

09:22 Garofalo sequence 9: TP_A_9.ob, TP_B_9.ob, TP_C_9.ob, TP_D_9.ob, TP_E_9.ob, FIELD_9.ob
  • 09:35 seeing 1.9"
  • 09:41 seeing 2.0"

09:48 Garofalo sequence 10: TP_A_10.ob, TP_B_10.ob, TP_C_10.ob, TP_D_10.ob, TP_E_10.ob, FIELD_10.ob
  • 09:50 seeing 2"
  • 09:53 seeing 3-4"
  • 10:05 seeing 5-6"

*10:09 Switch to MODS

10:15 CLOSE due to WIND GUST *weather loss

* LUCI DARKS:
  • 5 x 300 DIT, MER
  • 6 x 200 DIT, MER
  • 7 x 80 DIT, MER

-- %USERSIG{FeliceCusano - 2016-11-07}%

Comments

 
Topic revision: r9 - 03 Dec 2016, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback